# Quantum error correction for long chains of trapped ions

### v2

@@ V Conclusion
In future work, it would be interesting to explore the performance of other classes of small quantum LPDC codes  [60, 61, 62, 63, 64, 65] under the ion chain model.
VI Acknowledgments
@@ Appendix A Review of surface codes and BB codes
The distance-$d$ surface code is a $[[d^{2},1,d]]$ stabilizer code defined on a $d\times d$ two-colored square tiling as shown in Fig. 3. Qubits are placed on the $d^{2}$ vertices and each tile defines a stabilizer generator acting on the four incident qubits either as $X$ or $Z$ depending on the tile color. The stabilizer generators associated with boundary tiles act on two qubits only.
@@
To introduce BB6 codes, we define the matrices $x=Q_{\ell}\otimes I_{m}$ and $y=I_{\ell}\otimes Q_{m}$ where $m$ and $\ell$ are two positive integers. A BB6 code is defined by the pair of matrices
${\bf{H}}_{X}=[A_{1}+A_{2}+A_{3}~{}|~{}A_{4}+A_{5}+A_{6}],$
${\bf{H}}_{Z}=[A_{4}^{T}+A_{5}^{T}+A_{6}^{T}~{}|~{}A_{1}^{T}+A_{2}^{T}+A_{3}^{T}],$
@@ Appendix B Validation of the syndrome extraction tuning protocol
To validate the effectiveness of the syndrome extraction tuning protocol described in Algorithm 2, we execute this protocol to optimize the number of ancilla qubits $n_{a}$ of two BB6 codes with parameters $[[30,4,4]]$ and $[[48,4,6]]$ and we check that it provides a sensible choice.
Figure 4: Algorithm 2 chooses $n_{a}=3$ ancillas for $[[30,4,4]]$ BB6 code and $n_{a}=6$ for $[[48,4,6]]$ BB6 code.
To confirm that, we estimate the logical error rate of the two codes across a range of ancilla numbers $n_{a}=1,2,3,4,5,10,20$ as shown in Fig. 4. Algorithm 2 selects $n_{a}=3$ for the $[[30,4,4]]$ BB6 code and $n_{a}=6$ for the $[[48,4,6]]$ BB6 code. Fig. 4 clearly illustrates that increasing $n_{a}$ beyond these chosen values provides progressively smaller reductions in logical error rates, indicating diminishing returns. Specifically, for the $[[30,4,4]]$ BB6 code, using $n_{a}=20$ actually degrades the logical error rate compared to the protocol’s selection of $n_{a}=3$.
@@ Appendix C Fitting formulas for surface codes and BB5 codes
Most of the code simulations in previous papers employ a circuit model with parallel gate operations and uniform noise rates [2, 46, 47, 48]. We will refer to such a circuit model as the parallel circuit model. Fitting formulas for logical error rates of important code families such as surface codes and BB6 codes were studied under the parallel circuit model [2, 66, 47]. For surface codes, [2] presented a remarkably simple fitting formula
@@
where $d$ is the distance of the code, $p_{L}$ is the logical error rate, $p$ is the physical error rate, $c$ and $p_{\text{th}}$ are constants. For the parallel circuit model, the constants are $c=0.03$ and $p_{\text{th}}=0.0057$ [2]. Note that the same formula (1) with the same constants $c$ and $p_{\text{th}}$ effectively approximates the logical error rate of all surface codes. In contrast, [47] uses a more complicated formula $p_{L}=p^{d/2}e^{c_{0}+c_{1}p+c_{2}p^{2}}$ to fit BB6 codes, and each BB6 code instance requires a distinct set of constants $c_{0},c_{1},c_{2}$.
Under the ion chain model, we find that the formula (1) with constants $c=0.003$ and $p_{\text{th}}=0.0032$ provides a good fit for the 3 surface codes in Table 1. However, this formula does not generalize to longer surface codes with distance larger than 7. For BB5 codes, we use a more complicated fitting formula $p_{L}=p^{(d+1)/2}e^{c_{0}+c_{1}p+c_{2}p^{2}}$, where each BB5 code instance has its own set of constants $c_{0},c_{1},c_{2}$. These constants for the two BB5 code instances are listed in Table 3.

### v3

@@ 5 Conclusion
In future work, it would be interesting to explore the performance of other classes of small quantum LPDC codes [scruby2024high, lin2024quantum, wang2024coprime, eberhardt2024pruning, lin2025single, guemard2025moderate] under the ion chain model.
For simplicity, we assume that all the single-qubit operations share the same noise rate in the ion chain model. However, in practice, mid-circuit measurements might be noisier than unitary gates [pogorelov2025experimental]. Noise also varies over qubits and time. In future work, it would be interesting to explore non-uniform noise models including such features. We can start with a relatively simple observation. If we assign a higher error rate for measurements in our noise model, the optimal number of ancilla qubits $n_{a}$ will likely increase. This is because the number of measurements is given by $\lceil r/n_{a}\rceil$ (where $r$ is the number of syndromes) in Algorithm˜1. Noisier measurements make it beneficial to decrease the total number of measurements, which is achieved by increasing $n_{a}$.
For simplicity, we assume that all operations take the same amount of time expect measurements. The assumption that measurements are 30 times slower than two-qubit gates is realistic but single-qubit unitary gates are typically much faster [pogorelov2025experimental]. A future direction is to explore the performance of our quantum error correction scheme when assuming faster single-qubit gates.
It has been observed in [pogorelov2025experimental] that dephasing is the dominant error source on idling qubits in trapped ion hardware, meaning the probability of a $Z$ error is significantly higher than $X$ or $Y$ errors. For the sake of simplicity, our current model uses depolarizing errors for idling qubits, assigning equal probability to $X$, $Y$, and $Z$ errors. We have therefore deferred the incorporation of this physically observed biased noise into our ion chain model for future work. The XZZX variant of surface codes has been shown to perform significantly better than standard surface codes under such biased models [bonilla2021xzzx], making it an interesting open problem to modify BB5 and/or BB6 codes to similarly enhance their performance under biased noise.
Hadamard gates and controlled-Pauli gates used in Algorithm˜1 are typically non-native gates on a trapped-ion processor. The decomposition of non-native gates (such as controlled-Pauli and Hadamard gates) into native Mølmer–Sørensen and single-qubit operations, along with subsequent hardware-level optimizations (including gate elimination and commutation), is a substantial field of study in itself [sorensen1999quantum, sorensen2000entanglement, kreppel2023quantum]. To maintain the paper’s primary focus on the ion chain model and the new code construction, we have deferred this complex circuit optimization step to future work. Detailed analysis in this area will be necessary for achieving the lowest possible error rates during hardware implementation.
6 Acknowledgments
@@ Appendix A Review of surface codes and BB codes
The distance-$d$ surface code [^1] The version of surface code presented in Appendix A is usually referred to as the rotated surface code in the literature. We use the name surface code throughout the paper for simplicity. is a $d^{2},1,d]]$ stabilizer code defined on a $d\times d$ two-colored square tiling as shown in Fig.˜3. Qubits are placed on the $d^{2}$ vertices and each tile defines a stabilizer generator acting on the four incident qubits either as $X$ or $Z$ depending on the tile color. The stabilizer generators associated with boundary tiles act on two qubits only.
@@
Recall from Section˜4 that $Q_{\ell}$ is defined as the $\ell\times\ell$ circulant matrix whose first row is $(010\dots 0)$. To introduce BB6 codes, we define the matrices $x=Q_{\ell}\otimes I_{m}$ and $y=I_{\ell}\otimes Q_{m}$ where $m$ and $\ell$ are two positive integers. A BB6 code is defined by the pair of matrices
${\bf{H}}_{X}=[A_{1}+A_{2}+A_{3}\penalty 10000\ |\penalty 10000\ A_{4}+A_{5}+A_{6}],$
${\bf{H}}_{Z}=[A_{4}^{T}+A_{5}^{T}+A_{6}^{T}\penalty 10000\ |\penalty 10000\ A_{1}^{T}+A_{2}^{T}+A_{3}^{T}],$
@@ Appendix B Validation of the syndrome extraction tuning protocol
To validate the effectiveness of the syndrome extraction tuning protocol described in Algorithm˜2, we execute this protocol to optimize the number of ancilla qubits $n_{a}$ of two BB6 codes with parameters $[[30,4,4]]$ and $[[48,4,6]]$ and we check that it provides a sensible choice.
Figure 4: Algorithm˜2 chooses $n_{a}=3$ ancillas for $[[30,4,4]]$ BB6 code and $n_{a}=6$ for $[[48,4,6]]$ BB6 code.
To confirm that, we estimate the logical error rate of the two codes across a range of ancilla numbers $n_{a}=1,2,3,4,5,10,20$ as shown in Fig.˜4. Algorithm˜2 selects $n_{a}=3$ for the $[[30,4,4]]$ BB6 code and $n_{a}=6$ for the $[[48,4,6]]$ BB6 code. Fig.˜4 clearly illustrates that increasing $n_{a}$ beyond these chosen values provides progressively smaller reductions in logical error rates, indicating diminishing returns.
@@ Appendix C Fitting formulas for surface codes and BB5 codes
Most of the code simulations in previous papers employ a circuit model with parallel gate operations and uniform noise rates [fowler2012surface, tremblay2022constant, bravyi2024high, berthusen2025toward]. We will refer to such a circuit model as the parallel circuit model. Fitting formulas for logical error rates of important code families such as surface codes and BB6 codes were studied under the parallel circuit model [fowler2012surface, bravyi2013simulation, bravyi2024high]. For surface codes, [fowler2012surface] presented a remarkably simple fitting formula
@@
where $d$ is the distance of the code, $p_{L}$ is the logical error rate, $p$ is the physical error rate, $c$ and $p_{\text{th}}$ are constants. For the parallel circuit model, the constants are $c=0.03$ and $p_{\text{th}}=0.0057$ [fowler2012surface]. Note that the same formula (1) with the same constants $c$ and $p_{\text{th}}$ effectively approximates the logical error rate of all surface codes. In contrast, [bravyi2024high] uses a more complicated formula $p_{L}=p^{d/2}e^{c_{0}+c_{1}p+c_{2}p^{2}}$ to fit BB6 codes, and each BB6 code instance requires a distinct set of constants $c_{0},c_{1},c_{2}$.
Under the ion chain model, we find that the formula (1) with constants $c=0.003$ and $p_{\text{th}}=0.0032$ provides a good fit for the 3 surface codes in Table˜1. However, this formula does not generalize to longer surface codes with distance larger than 7. For BB5 codes, we use a more complicated fitting formula $p_{L}=p^{(d+1)/2}e^{c_{0}+c_{1}p+c_{2}p^{2}}$, where each BB5 code instance has its own set of constants $c_{0},c_{1},c_{2}$. These constants for the two BB5 code instances are listed in Table˜3.